Solve the following equation for x:
6(x+6)-8x=28
Simplify your answer as much as possible.

Answer by stanbon(75887)

Solve the following equation for x:
6(x+6)-8x=28
6x+36-8x = 28
-2x = -8
x = 4
========
Cheers,
Stan H.